# Adiabatic Fine-Tuning of Neural Quantum States Enables Detection of Phase Transitions in Weight Space

### v1

@@ 1 Introduction and Related Work
The simulation of quantum many-body systems represents one of the most challenging problems in computational physics, primarily due to the exponential growth of the Hilbert space with system size (Feynman, 2018; Nielsen & Chuang, 2010). Neural Quantum States (NQS) have emerged as a promising approach, leveraging the expressive power of deep learning to represent quantum wavefunctions (Carleo & Troyer, 2017; Medvidović & Moreno, 2024). In this approach, neural networks serve as ansätze for the wavefunction, mapping quantum states to their corresponding probability amplitudes. While these models give very good results in ground state energy calculations (Chen & Heyl, 2023; Valenti et al., 2022; Wu et al., 2024; Rende et al., 2024), it remains challenging to understand how they encode physical information.
Our work explores the relationship between phase transitions and the geometry of neural network weight spaces. In deep learning, various studies have explored the structure of weight space with diverse objectives, ranging from analyzing the loss landscape to understand mode connectivity (Garipov et al., 2018; Entezari et al., 2021), to leveraging neural network weights as data for downstream predictions (Eilertsen et al., 2020; Unterthiner et al., 2020; Schürholt et al., 2024). Meanwhile, in quantum physics, phase transitions are traditionally characterized through order parameters and correlation functions. The intersection of these fields has led to emerging research on how neural networks capture physical symmetries and order parameters (Wetzel et al., 2020; Frohnert & van Nieuwenburg, 2024; Cybiński et al., 2024), and how it’s possible to exploit machine learning models to learn about physical phase transitions (Van Nieuwenburg et al., 2017; Dawid et al., 2020; Greplova et al., 2020; Arnold et al., 2021). However, directly analyzing the weights of neural quantum states as a means of detecting phase transitions remains largely unexplored.

### v2

@@ 1 Introduction and Related Work
The simulation of quantum many-body systems represents one of the most challenging problems in computational physics, primarily due to the exponential growth of the Hilbert space with system size (Feynman, 2018; Nielsen & Chuang, 2010). Neural Quantum States (NQS) have emerged as a promising approach, leveraging the expressive power of deep learning to represent quantum wavefunctions (Carleo & Troyer, 2017; Medvidović & Moreno, 2024). In this approach, neural networks serve as ansätze for the wavefunction, mapping quantum states to their corresponding probability amplitudes. While these models give very good results in ground state energy calculations (Chen & Heyl, 2023; Valenti et al., 2022; Wu et al., 2024; Rende et al., 2024b), it remains challenging to understand how they encode physical information.
Recent work by Rende et al. (Rende et al., 2024a) introduced a fine-tuning strategy in which a network pretrained near a phase transition is adapted across a phase diagram by updating only the output layer, reducing computational costs while maintaining accuracy across phases. Building on this idea, we investigate how phase transitions relate to the geometry of neural network weight space. In deep learning, various studies have explored the structure of weight space with diverse objectives, ranging from analyzing the loss landscape to understand mode connectivity (Garipov et al., 2018; Entezari et al., 2021), to leveraging neural network weights as data for downstream predictions (Eilertsen et al., 2020; Unterthiner et al., 2020; Schürholt et al., 2024). Meanwhile, in quantum physics, phase transitions are traditionally characterized through order parameters and correlation functions. The intersection of these fields has led to emerging research on how neural networks capture physical symmetries and order parameters (Wetzel et al., 2020; Frohnert & van Nieuwenburg, 2024; Cybiński et al., 2024), and how it’s possible to exploit machine learning models to learn about physical phase transitions (Van Nieuwenburg et al., 2017; Dawid et al., 2020; Greplova et al., 2020; Arnold et al., 2021). However, directly analyzing the weights of neural quantum states as a means of detecting phase transitions remains largely unexplored.